THE ELECTRIC CURRENT I, IN AMPERES, IN A CIRCUIT VARIES DIRECTLY AS THE VOLTAGE V. WHEN 12 VOLTS ARE APPLIED, THE CURRENT IS 4 AMPERES. WHAT IS THE CURRENT WHEN 18 VOLTS ARE APLLIED?
Answer by stanbon(75887)

THE ELECTRIC CURRENT I, IN AMPERES, IN A CIRCUIT VARIES DIRECTLY AS THE VOLTAGE V. WHEN 12 VOLTS ARE APPLIED, THE CURRENT IS 4 AMPERES. WHAT IS THE CURRENT WHEN 18 VOLTS ARE APPLIED?
-----------
I = kV
Find "k":
4 = k*12
k = 4/12 = 1/3
-----------
Equation: I = (1/3)V
If 18 volts are applied you get:
I = (1/3)*18
I = 6 amps.
==============
Cheers,
Stan H.